# Quality Time: Carbon-Aware Quality Adaptation for Energy-Intensive Services

### v1

@@ 6. Discussion
We highlight the future impact of carbon efficiency in energy-intensive services and line out the limitations of this study.
Impact
@@
We want to emphasize the extend of carbon efficiency improvements for cloud services, based on our analysis in Section 4.3. ChatGPT, one of the most energy-intensive cloud services today, reached over 200 million active users in 2024 (Reuters, 2024), with fewer than 10 million paying customers (Tarantola, 2024). Assuming just 5 requests per user per day, this amounts to nearly 42 million requests per hour on average, which is an order of magnitude more than what we consider in our scenarios. Assuming performance similar to Llama 3.1 70B, we estimate that the annual operational carbon emissions of large-scale LLM service providers currently exceed multiple 10,000 tCO2—a figure set to rise with multi-modal AI models.
@@
Furthermore, this study assumes ACI as the most likely metric to be implemented in official policy. However, and increasing part of the community is advocating to base load shifting decisions based on marginal data (WattTime, 2022; Buchanan et al., 2023; Dodge et al., 2022), which—if implemented properly—better reflects the underlying complexities of electric grids. However, to this day, existing methodologies for MCI are based on predicting the marginal generator, which is ambiguous and error prone. Moreover, current reporting lacks spatial granularity to capture effects like energy curtailments from local grid congestion. We predict, that with improved methodologies and more granular MCI reporting, carbon-aware QoR adaptation could have much greater potential, as MCI shows higher temporal variability than ACI, enabling the use of shorter validity periods.
Limitations
In this paper, we provide first insights into the potential of carbon-aware QoR adaptation based on an LLM inference use case. Applying our approach to more complex, real-world scenarios with multiple user groups, service quality tiers, and heterogeneous infrastructure is possible from a modeling perspective, but not yet investigated. Furthermore, we only concentrate on QoR within a single data center and do not distribute load across multiple geo-distributed regions—which is known to yield the largest potential for carbon savings (Sukprasert et al., 2024b). This paper represents the first comprehensive analysis of QoR adaptation for energy and carbon efficiency, and sets the stage for more detailed explorations in future research.
@@
Second, our current modeling only considers the QoR for user groups, rather than individual users. This assumes a uniform distribution of users within each group and may lead to fairness concerns if the distribution is uneven. For example, ACI tends to peak in the evenings in most regions, meaning that users who usually access the service during these hours can experience a lower QoR than other users in their group. We consider addressing fairness aspects with more granular user modeling as as an area for future work.
@@ 7. Related Work
Carbon-aware computing
The large majority of approaches in carbon-aware computing focuses on shifting flexible batch processing towards times (Wiesner et al., 2021; Radovanovic et al., 2022; Fridgen et al., 2021; Lechowicz et al., 2023) and locations (Zheng et al., 2020; Lindberg et al., 2022; Wiesner et al., 2024) with clean energy. Hence, for batch jobs, the limitations and trade-offs of carbon-aware spatio-temporal workload shifting have been investigated thoroughly (Hanafy et al., 2024; Sukprasert et al., 2024b). However, carbon efficiency in continuously running services that usually have little to no delay-tolerance has received only little attention so far.
Existing approaches for carbon-aware services primarily focus on load balancing across geo-distributed datacenters and the respective capacity right-sizing under SLA constraints (Zhou et al., 2013, 2016). For example, Casper (Souza et al., 2023) is a carbon-aware scheduling and provisioning system for web services that minimizes the carbon footprint while respecting latency constraints. CDN-Shifter (Murillo et al., 2024) explores carbon-aware spatial load shifting for content delivery networks. Caribou (Gsteiger et al., 2024) offloads serverless workflows across geo-distributed regions based on average carbon intensity. However, no existing approaches address services that must operate in specific data centers due to data, security, or infrastructure constraints.
Sustainability of cloud services
Although the carbon footprint of AI training has been an active topic of research for some time (Patterson et al., 2021; Dodge et al., 2022; Strubell et al., 2020; Dhar, 2020) only recently there has been an increasing focus on AI inference (Chien et al., 2023; Eilam et al., 2023). Especially the rapid adoption of LLMs has resulted in multiple studies that try to quantify their growing carbon footprint (Nguyen et al., 2024; Faiz et al., 2024; Luccioni et al., 2024).
While service degradation is a common practice to prevent overload (Meza et al., 2023; Park et al., 2024), it has already been proposed to design applications to offer different quality tiers of a service for sustainability (Vitali, 2022). In the context of carbon-aware LLMs, there only exists a single approach, Sprout (Li et al., 2024b), which optimizes an LMMs autoregressive generation process. In particular, the authors propose to return shorter responses during periods of high carbon intensity, as the energy demand of a response directly correlates with the amount of generated tokens. Sprout clearly classifies as an approach which adapts QoR over time and can benefit from our scheduling decisions.
@@ 8. Conclusion
As data centers and cloud computing infrastructure continue to expand, their role in emissions reduction will be crucial for meeting global climate targets. In this paper, we explored the potential of adjusting a cloud service’s QoR to improve both energy and carbon efficiency. Our findings demonstrate that, in many regions, carbon-aware QoR adaptations can reduce emissions by up to 10 %, beyond the savings from reduced energy demand alone. Moreover, we propose an approach for dynamically adjusting the QoR of best-effort users to stay within a predefined annual carbon budget, which is an important contribution for service providers facing increasing pressure from carbon reporting and pricing.
@@ Acknowledgements.
Appendix A CarbonCast MAPE
@@
Table 2. Carbon forecast MAPE over 96 hours for all observed regions as reported in CarbonCast (Maji et al., 2023)

### v2

@@ 6. Discussion
This section highlights the impact of our results on energy-intensive services and discusses the limitations of this study.
@@
Impact.  We emphasize the significant potential for carbon efficiency improvements in cloud services, based on our analysis in Section 5.2. ChatGPT, one of the most energy-intensive cloud services today, reached over 200 million active users in 2024 (Reuters, 2024). Assuming just 5 requests per user per day, this amounts to nearly 42 million requests per hour on average, which is an order of magnitude more than what we consider in our larger scenarios. Assuming performance similar to Llama 3.1 70B, we estimate that the annual operational carbon emissions of large-scale LLM service providers currently exceed multiple 10,000 tCO2—a figure set to rise with multi-modal AI models.
@@
Furthermore, this study assumes ACI as the most likely metric to be implemented in official policy. However, and increasing part of the community is advocating to base load shifting decisions based on marginal signals (WattTime, 2022; Buchanan et al., 2023; Dodge et al., 2022), which—if implemented properly—better reflects the underlying complexities of electric grids. Yet, to this day, existing methodologies for MCI are based on predicting the marginal generator, which is ambiguous and error prone. Moreover, current reporting lacks spatial granularity to capture effects like energy curtailments from local grid congestion. We predict that with improved methodologies and more granular MCI reporting, carbon-aware QoR adaptation could have much greater potential, as MCI shows higher temporal variability than ACI, enabling the use of shorter validity periods. Our approach is directly applicable to MCI signals or any other scalar metric without requiring adaptations.
Limitations and future work.  This paper presents the first analysis on the potential of carbon-aware QoR adaptation in services, focusing on an LLM inference use case. While our approach is already applicable to more complex scenarios with multiple service quality tiers and heterogeneous infrastructure, such cases have not yet been investigated experimentally.
@@
Second, our current modeling only considers the QoR for user groups, rather than individual users. This assumes a uniform distribution of users within each group and may lead to fairness concerns if the distribution is uneven. For example, ACI tends to peak in the evenings in most regions, meaning that users who usually access the service during these hours can experience a lower QoR than other users in their group. We consider fairness aspects with more granular user modeling as an area for future work.
Third, future work should explore the impact of adaptive QoR adjustments on user quality of experience and the resulting implications on user behavior. For instance, in the context of generative AI, reduced QoR might incentivize users to resubmit requests, potentially offsetting carbon savings. Unfortunately, to the best of our knowledge, there are no established behavioral models for generative AI usage. Developing such models would be valuable to better quantify the real-life impacts of sustainability approaches, including rebound effects.
@@ 7. Related Work
Carbon-aware cloud services.  The large majority of approaches in carbon-aware computing focuses on shifting flexible batch processing towards times (Wiesner et al., 2021; Radovanovic et al., 2022; Fridgen et al., 2021; Lechowicz et al., 2023) and locations (Zheng et al., 2020; Lindberg et al., 2022; Wiesner et al., 2024) with clean energy. Hence, for batch jobs, the limitations and trade-offs of carbon-aware spatio-temporal workload shifting have been investigated thoroughly (Hanafy et al., 2024b; Sukprasert et al., 2024b; Hanafy et al., 2024a; Lechowicz et al., 2024). However, carbon efficiency in continuously running services that usually have little to no delay-tolerance has received only little attention so far.
Existing approaches for carbon-aware services primarily focus on load balancing across geo-distributed data centers and the respective capacity right-sizing under SLA constraints (Zhou et al., 2013, 2016). For example, Casper (Souza et al., 2023) is a carbon-aware scheduling and provisioning system for web services that minimizes the carbon footprint under latency constraints. CDN-Shifter (Murillo et al., 2024) explores carbon-aware spatial load shifting for content delivery networks. Caribou (Gsteiger et al., 2024) offloads serverless workflows across geo-distributed regions based on average carbon intensity. However, there are no existing approaches that address services which must operate in a single region due to data, security, or infrastructure constraints.
Sustainability in generative AI inference.  Although the carbon footprint of AI training has been an active topic of research for some time (Patterson et al., 2021; Dodge et al., 2022; Strubell et al., 2020; Dhar, 2020) only recently there has been an increasing focus on AI inference (Chien et al., 2023; Eilam et al., 2023; Samsi et al., 2023). Especially, the rapid adoption of LLMs has resulted in multiple studies that try to quantify their growing carbon footprint (Nguyen et al., 2024; Faiz et al., 2024; Luccioni et al., 2024). Many recent approaches target sustainability aspects of inference services throuh power-aware scheduling (Stojkovic et al., 2024, 2025; Qiu et al., 2025; Patel et al., 2024) or by considering their carbon and water footprint in geographic load balancing (Li et al., 2024d).
Service quality adaptation.  While service degradation is a common strategy to prevent system overload (Meza et al., 2023; Park et al., 2024), recent research has proposed designing applications with different service quality tiers as a means of improving sustainability (Vitali, 2022). In the context of carbon-aware LLMs, we identified only one approach, called Sprout (Li et al., 2024b), which optimizes the autoregressive generation process. Sprout returns shorter responses during periods of high carbon intensity, as the energy demand directly correlates with the number of generated tokens. Thereby, it effectively adapts QoR over time and can be further enhanced by leveraging the optimized infrastructure deployments and load balancing proposed in this paper.
@@ 8. Conclusion
As data centers and cloud computing infrastructure continue to expand, their role in emissions reduction will be crucial for meeting global climate targets. In this paper, we explored the potential of adjusting a cloud service’s QoR to improve both energy and carbon efficiency. Our findings demonstrate that, in many regions, carbon-aware QoR adaptations can reduce emissions by up to 10 %, beyond the savings from reduced energy demand. Moreover, we propose an approach to automatically adapt the QoR of best-effort users to stay within a predefined annual carbon budget, which is an important contribution for service providers facing increasing pressure from carbon reporting and pricing.
@@ Acknowledgements.
Appendix A Proof of Theorem 3.1
Theorem 3.1 We assume $p_{m,q}^{\text{max}}>p_{m}^{\text{idle}}>0$ and $n\in[0,1]$. When minimizing (10) subject to (11) – (14) from the perspective of data center operators, the utilization-dependent power model,
$p^{i}_{m,q}=p_{m}^{\text{idle}}+\bigl{(}p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\bigl{)}\cdot\bigl{(}\text{util}^{i}_{q}\bigl{)}^{n},$
can be simplified to
$p^{i}_{m,q}=p_{m,q}^{\text{max}},$
as it always yields an equivalent optimal deployment $D^{i}$.
Lemma A.1.
Increasing the utilization of machines does not result in additional penalties, so there is no incentive to deploy more machines than strictly necessary. Formally, if $p_{m,q}^{\text{max}}>p_{m}^{\text{idle}}>0$ and $n\in[0,1]$, an utilization-dependent power model $f$ is strictly subadditive
(18) $f(x+y)<f(x)+f(y),$
where $x$ and $y$ represent utilization levels $\text{util}^{i}_{q}\in(0,1]$.
Proof of Lemma A.1.
We show that $f$ is strictly subadditive for the two cases $n=1$ and $n<1$.
Case 1: Linear power model ($n=1$):
(19) $f(x+y)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot\left(x+y\right)$
$f(x+y)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)x$
(20) $\quad+p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)y-p_{m}^{\text{idle}}$
(21) $f(x+y)$ $=f(x)+f(y)-p_{m}^{\text{idle}}$
(22) $f(x+y)$ $<f(x)+f(y)$
Case 2: Sub-linear power model ($n<1$): For a strictly concave function $f$ with $f(0)\geq 0$, $f$ is strictly subadditive on $(0,\infty)$.
(23) $f(x)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot x^{n}$
(24) $f^{\prime}(x)$ $=n\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot x^{n-1}$
(25) $f^{\prime\prime}(x)$ $=\underbrace{n(n-1)\vphantom{\Bigl{(}}}_{<1}\ \underbrace{\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)}_{>1}\cdot\underbrace{x^{n-2}\vphantom{\Bigl{(}}}_{>1}<0$
(26) $f^{\prime\prime}(x)<0$ $\implies f\text{ is strictly concave}$
Additionally, $f(0)=p_{m,q}^{\text{max}}>0$. ∎
Proof of Theorem 3.1.
For each interval $i$, we define the minimal deployment $\check{D}^{i}$ as any feasible solution that uses the smallest possible number of machines:
[…]
  1. (1)
    Feasibility with fewer machines is impossible. If a solution $\widehat{D}^{i}$ has fewer machines than $\check{D}^{i}$ in at least one dimension, it would violate Sufficient resources (12).
  2. (2)
    Using extra machines is suboptimal. By Lemma A.1, under the utilization-dependent power model, there is no incentive to deploy extra machines. Likewise, any extra machine increases the objective under the simplified power model.
We conclude that, for both power models, any optimal solution must yield an equivalent deployment $\check{D}^{i}$ for each interval $i$. ∎
Appendix B Request Traces
This section describes all request trace datasets used in the experiments. The traces are visualized in Figure 12; Table 3 presents summary statistics.
Table 3. Request trace and 24-hour forecast statistics.
Dataset statistics ($1\times 10^{6}$) Forecast statistics
Mean ± Std Dev Min Max MAPE
Static 1.00 ± 0.00 1.00 1.00 0.0 ± 0.0
Normal 1.00 ± 0.34 0.00 2.36 38.6 ± 24.6
Wiki (en) 3.38 ± 0.80 1.88 16.41 13.9 ± 8.4
Wiki (de) 0.42 ± 0.24 0.04 1.56 32.1 ± 15.3
Taxi 0.33 ± 0.14 0.04 0.71 26.5 ± 7.1
Cell B 1.94 ± 0.61 0.73 4.10 27.2 ± 13.5
Cell D 2.87 ± 0.80 1.02 7.76 22.1 ± 15.8
Cell F 1.58 ± 0.41 0.87 4.32 18.2 ± 9.3
Figure 12. First month of data and KDE over all data points in 2023 for the eight datasets used in the evaluation.
Artificial Datasets.  We artificially create two traces:
  • Static assumes a constant stream of $1\times 10^{6}$ hourly requests.
  • Random comprises points randomly sampled from a normal distribution with a mean of $\gamma=1\times 10^{6}$ and a standard deviation of $\sigma=0.33\times 10^{6}$.”
During online optimization, we always assume the mean of $1\times 10^{6}$ requests per hour as forecasts.
Real Datasets.  We create three traces based on real data:
  • Wiki (en) is based on the Wikipedia pageview statistics (Wikimedia Foundation, [n. d.]), namely all hourly requests to en.wikipedia.org.
  • Wiki (de) represents all requests to de.wikipedia.org. The trace differs notably from Wiki (en), as the German Wikipedia is primarily accessed from a single timezone, while the English Wikipedia sees traffic from around the globe.
  • Taxi is based on hourly aggregates of taxi trips in New York City (Taxi and Commission, 2025). We sum over all trip records (Yellow Taxi, Green Taxi, For-Hire Vehicle, High Volume For-Hire Vehicle) and multiply the events by factor 10, to bring them into a similar range with the other datasets.
To simulate realistic forecasts, we are using using Prophet (Taylor and Letham, 2017), a state-of-the-art forecasting model that fits daily, weekly, and annual seasonalities. In particular, we fit a model every day at midnight using 3 years of historical data to forecast the rest of the year. The predictions exhibit realistic errors, denoted in Table 3.
Synthetic Datasets.  As all real datasets exhibit very periodic patterns, we additionally created three synthetic datasets to demonstrate that carbon-aware QoR adaptation is feasible on highly unpredictable traces. For this, we aggregated the instance events of the Google cluster traces (version 3) (Google, 2019) individually for the eight available Borg cells. Data was retrieved via the GoogleSQL query:
SELECT
  TIMESTAMP_TRUNC(TIMESTAMP_MICROS(time), HOUR) AS hour,
  COUNT(*) AS num_events
FROM ‘google.com:google-cluster-data.<TABLE>‘
-- Exclude invalid values
WHERE time > 0 AND time < 9223372036854775807
GROUP BY hour
ORDER BY hour;
[…]
We identified the three traces with the lowest 24-hour autocorrelation coefficients, i.e. the traces that exhibit the least daily seasonality: Cell B (0.17), Cell D (0.27), and Cell F (0.22). Since the Google cluster dataset only covers one month of data, we fit a DoppelGANger (Lin et al., 2019) model for each of the traces, which is a state-of-the-art GAN for timeseries generation, and generate 4 years of synthetic data per cell. We apply the same forecasting methodology that is used for real datasets.
Appendix C Carbon Intensity Forecasts
For long-term carbon intensity forecasts, we apply the same forecasting methodology that is used for the request trace datasets. As in (Lechowicz et al., 2024), we generate synthetic short-term forecasts of up to 4 days by adding Gaussian noise to the actually observed carbon intensity data to match the MAPEs reported by the state-of-the-art forecasting model CarbonCast (Maji et al., 2023) for each region, see Table 4. Short-term forecasts are updated daily at midnight.
Table 4. Carbon forecast MAPE over 96 hours as reported in CarbonCast (Maji et al., 2023)